OpenStudy (anonymous):

heellpppp..... solve using substitution u=x+3 2(x+3)^2-17(x+3)=-21

OpenStudy (anonymous):

\[2u^2-17u+21=0\] is a start

OpenStudy (anonymous):

by some miracle this one factors as \[(2u-3)(u-7)=0\] and so you can solve for \(u\) easily, then replace \(u\) by \(x+3\) and solve for \(x\)

OpenStudy (anonymous):

ok so before i plug in back the x+3 i got u=7 and for my other solution it is negative 3/2? right?

OpenStudy (anonymous):

never mind its positive

OpenStudy (anonymous):

its positive but then you need to use 3/2-3. Then solve. Remember to use the same denom.

OpenStudy (anonymous):

You should get the right answer. Then plug it back in to see if you equation is right.

OpenStudy (anonymous):

Could you explain how did you get the factors satellite73?